Search results for If v=(y²+z²-x²) î+ (z²+x²-y²) j^+(x²+y²-z²) k^, then find to it's divergence v and curl v? 



1
Expert's answer
2021-12-15T08:12:29-0500


div "\overrightarrow{V}=\frac{\partial}{\partial\>x}(y^2+z^2-x^2)+\frac{\partial}{\partial\>y}(z^2+x^2-y^2)+\frac{\partial}{\partial\>z}(x^2+y^2-z^2)"


"=-2-2y-2z"

"=-2(x+y+z)"




"Curl{\overrightarrow{V}}" ."\begin{vmatrix}
 i & j&k\\
 \frac{\partial}{\partial\>x}&\frac{\partial}{\partial\>y}&\frac{\partial}{\partial\>z} \\
y^2+z^2-x^2&z^2+x^2-y^2&x^2+y^2-z^2
\end{vmatrix}"



"=(2y-2z)\utilde{i}+(2z-2x)\utilde{j}+(2x-2y)\utilde{k}"


"=2(y-z)\utilde{i}-2(x-z)\utilde{j}+2(x-y)\utilde{k}"